An object accelerates at point O. At a point 5m from point O the object hits a velocity of 3m/s-1 if the acceleration is 4m/s-2


Find the object's position and velocity at t = 2s.


Determine the object's position when velocity is 5m/s-1

1
Expert's answer
2021-03-17T16:53:20-0400

Answer

Velocity of object at t=2 s

"v=4\times 2=8m/s"


Position

"x=8\times 2=16m"


By equation of motion

"v^2=u^2+2as"

Initially u=3m/s

So Displacement


"s=\frac{5^2-3^2}{2\times 4}=2m"

Position at this point x=5+2=7m